# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ Abstract
We introduce a general diploid population model with self-fertilization and possible overlapping generations, and study the genealogy of a sample of $n$ genes as the population size $N$ tends to infinity. Unlike traditional approach in coalescent theory which considers the unconditional (annealed) law of the gene genealogies averaged over the population pedigree, here we study the conditional (quenched) law of gene genealogies given the pedigree. We focus on the case of high selfing probability and obtain that this conditional law converges to a random probability measure, given by the random law of a system of coalescing random walks on an exchangeable fragmentation-coalescence process of Ber (04). This system contains the system of coalescing random walks on the ancestral recombination graph as a special case, and it sheds new light on the site-frequency spectrum (SFS) of genetic data by specifying how SFS depends on the pedigree. The convergence result is proved by means of a general characterization of weak convergence for random measures on the Skorokhod space with paths taking values in a locally compact Polish space.
@@ 1 Introduction
This conceptual flaw was not unrecognized BNA (90); WKLR (12); WKW (16); WBLW (17); Ral (19), but perhaps luckily, it turned out that the Kingman coalescent can still be applied under the standard assumptions of neutral coalescent model because then the conditional limit is still equal to the unconditional limit Tyu (15). However, when there are simultaneous multiple mergers, the conditional genealogy can be different from the unconditional genealogy DFBW (24).
@@
In ABFW (25), the authors study the diploid Cannings model introduced in BLS (18), where selfing is excluded. This is a model in which the offspring distribution, described by the matrix $(V_{ij})$ of the number of offspring between all pairs of parents, is invariant under permutations of the parents. Under the same condition on the offspring distribution that guaranteed annealed convergence to the $\Xi$-coalescent holds, the authors in ABFW (25) showed that the genealogies conditioned on the pedigree converge to an inhomogeneous $(\Psi,c_{\rm pair})$-coalescent, where $\Psi$ is a Poisson point process on $0,\infty)\times(\Delta\setminus\{\bf 0\})$ with intensity $dt\otimes\frac{1}{\langle x,x\rangle}\Xi(dx)$. This limiting process consists of the independent superposition of a Kingman coalescent with rate $c_{\rm pair}$ and multiple merger events whose the timing and intensity are specified by $\Psi$. The result in ABFW (25) generalizes earlier work in Tyu (15) and DFBW (24).
@@
While the Cannings model in BLS (18); ABFW (25) is quite general and captures a wide range of reproductive variance, it has two limitations from a biological perspective. Firstly, it does not allow for selfing (self-fertilization), an important evolutionary force found in taxa including eukaryotic microbes and marine invertebrates (SR, 17; YSH, 23) and is common in plants (AG, 89; HBG, 17; TWL+, 23). The empirical distribution of selfing probabilities among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ and some species reaching as high as 0.99 SL (85); AG (89); VK (01); BF (05); GKE (05); SAAMT (20). Secondly, the model assumes non-overlapping generations, a simplification that excludes biologically realistic scenarios in which individuals from different age cohorts may reproduce simultaneously. Overlapping generations are common in many natural populations and can introduce temporal correlations and ancestral dependencies not captured by discrete-generation models.
@@
The recent work NWF (25) introduces a diploid Moran model with a high selfing probability $\alpha_{N}$. Quenched coalescent limit for the coalescence time of a sample of size $n=2$ was obtained for that Moran model. Conditioning coalescence on the pedigree reveals three markedly different behaviors, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$.
@@
This paper aims to contribute to building the emerging quenched-coalescent theory, by going beyond the existing mathematical work Tyu (15); DFBW (24); NWF (25); ABFW (25). We introduce a general diploid model with overlapping generations that extends the haploid model of SW (08) and the diploid models of BLS (18) and NWF (25). Our model explicitly incorporates a selfing rate parameter and strictly contains the diploid Moran-type models with selfing in NWF (25); CLJ (22); Lin (09) as special cases. Our main result significantly strengthens the analysis in NWF (25), in the ‘limited outcrossing’ regime, by considering arbitrary sample size $n$ and a substantially broader class of models. Moreover, it complements the results of ABFW (25); Tyu (15); DFBW (24) by identifying a distinct limiting conditional coalescent process that arises in the high selfing regime.
@@
The new class of coalescent models described herein is a family of coalescing random walks on a directed random graph, which we call a $Q$-$\lambda$ graph in this paper. This random graph is a subgraph of an exchangeable fragmentation-coalescence process (EFC) introduced in Ber (04), where $Q$ is the rate matrix for coalescence and $\lambda$ is the fragmentation rate at which each node spits into two. The parameter $\lambda$ is the relative rate of outcrossing to coalescence as $N\to\infty$. In the regime where the classical (annealed) limit is a time-rescaled Kingman, we find that these $Q$-$\lambda$ graphs correspond to the ancestral recombination graphs (also known as ancestral selection graph) Hud (83); GM (97); KN (97) which are used to describe the genealogy of a sample of genome sequences with recombination.
The proof of the main result involves a novel characterization of convergence of random measures on the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for any locally compact Polish space $E$, and a general theorem about quenched convergence of Markov processes on a suitably enriched space of partitions. These results are then combined with a combinatorial argument to characterize the scaling limits of our diploid Sargasyan-Wakeley model. These results also allow us to strengthen the results of (ABFW, 25, Theorem 3.8) from quenched convergence in finite-dimensional distribution to weak convergence in distribution of random measures.
@@
From the application standpoint, the main motivation of coalescent theory (and our quenched-coalescent theory) is to describe the patterns of genetic variation expected under various biological scenarios to provide frameworks for statistical inference about past events and processes affecting populations. In a given population, gene genealogies are tree structures which emerge from tracing the ancestral lines of these sample backwards in time until the most recent common ancestor (MRCA). Mutations occurred in the past, along the ancestral lines, result in genetic diversity among the sample. For example, the site-frequency spectrum (SFS) is a commonly used measure of genetic variation upon which statistical inferences are based BWSH (01); Ach (09); EBBF (15); LF (15); GK (16); FLW+ (17); FKM+ (23). For a sample of $n$ genomes, the SFS records the number of polymorphic sites where a mutant base is found in $r\in\{1,\ldots,n-1\}$ copies Taj (89); BHK+ (95); Fu (95). Because per-site mutation rates are typically very small and the number of sites is large, the SFS is taken to reflect the total length of branches in the gene genealogy with $r$ descendants in the sample, or which are ancestral to $r$ of the samples. Different biological phenomena, such as population growth and natural selection, lead to different coalescent predictions for the SFS. Our main result sheds new light on the SFS of genetic data, by specifying how SFS depends on the pedigree. More precisely, under the conditional coalescent, the SFS should be viewed as a conditional SFS given the single pedigree as a latent variable. In Figure 5, we provide simulations for the conditional SFS given 5 different pedigrees and thus illustrate how the pedigree can impact the SFS of the data.
@@
Organization of this paper. Section 2 introduces our diploid exchangeable model with self-fertilization and overlapping generations. An annealed version of our main result, akin to classical results of coalescent theory such as those of Kin (82); Möh (98); Pit (99); Sag (99); Sch (00), is presented in Section 3. Section 4 presents the main quenched convergence result for the coalescent of our exchangeable model model conditional on the pedigree. In Section 5 we present applications of our main convergence result and illustrate how the SFS can depend strongly on the pedigree. The remainder of the paper focuses on the tools and calculations necessary to prove the main results: weak convergence of random measures in distribution in Section 6, and a suitable notion of convergence of the random pedigree and continuity of the coalescent law’s dependence thereon in Section 7.
@@ 2.1 The model
In this paper, we consider a diploid, monoecious, panmictic (well-mixed and randomly mating) population of constant size $N$, evolving in discrete time-steps with overlapping generations. This model generalizes the diploid Cannings model in ABFW (25); BLS (18) to allow overlapping generations and selfing, and extends the haploid model in SW (08) to diploid populations while supporting general offspring distributions.
@@
We consider discrete time-steps indexed by $k\in\mathbb{Z}_{+}=\{0,1,2,\ldots\}$, where $k=0$ is the present, $k=1$ the previous time-step, and so on backward into the past. Let $\big\{(K^{(k)}_{N},\textbf{V}^{(k)})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random variables that have the same distribution as $(K_{N},\textbf{V})$. Reproduction events in different time-steps are taken to be independent, and are described as follows:
For each $k\in\mathbb{Z}_{+}$, $K^{(k)}_{N}$ individuals are chosen uniformly without replacement, among the $N$ individuals from the $k$-th time-step, to be children. To describe parentage, we suppose that the individuals at each time-step are labeled by $[N]$, and we let $V^{(k)}_{i,j}$ be the number of offspring produced by the pair of individuals $(i,j)$ in the $(k+1)$-th time-step in the past. The $K_{N}^{(k)}$ individuals defined by the entries of $\textbf{V}^{(k)}$ constitute the new individuals in time-step $k$. Each individual has two sets of chromosomes, as it is a diploid population. Genetic lineages at an autosomal locus are transmitted according to Mendel’s law of random segregation, which means each gene copy in the offspring chooses independently from the two gene copies in the parent from which the gene copy is inherited.
@@
We further assume that each of the $K^{(k)}_{N}$ children independently chooses to have either a single parent (selfing) or two distinct parents (outcrossing), with probabilities $\alpha_{N}$ and $1-\alpha_{N}$ respectively. Therefore, conditional on $K_{N}^{(k)}$, the number of selfed offspring $S_{N}^{(k)}:=\sum_{i=1}^{N}V_{i,i}^{(k)}$ is an independent binomial random variable: $\mathrm{Bin}(K_{N}^{(k)},\alpha_{N})$. In other words, we assume that $(K_{N},\textbf{V})$ satisfies
@@
Figure 1: An illustration of our population model between time-steps $k+1$ and $k$ in the past, with size $N=6$. The number of children, whose edges are colored in black, is $K_{N}=3$. Further, $S_{N}=2$ as two of the children have reproduced by selfing. Individuals $1,2,4$, whose edges are marked in orange, simply persisted between consecutive time-steps and are not children. To demonstrate how to read the $V_{i}$ and $V_{i,i}$ from the pedigree, we see that $V_{1}=0$ as the first individual has no offspring. $V_{2}=2$ as it has the third and fourth individuals as offspring. $V_{2,2}=1$ as the third individual is a child via offspring while the fourth is via selfing.
@@
The population model described above includes two overlapping but distinct classes of models in the literature as special cases.
@@
When $K_{N}=N$ (non-overlapping generations) and $\alpha_{N}=0$ (no selfing), our model reduces to the Cannings model introduced in BLS (18) and studied in ABFW (25).
@@
Our model also generalizes the model in SW (08) in two ways: to incorporate more general offspring distribution, and from haploid to diploid.
Suppose the random symmetric matrices $\{V^{(k)}\}_{k\in\mathbb{Z}_{+}}$ satisfy the following extra assumption. Namely, their distribution is parametrized by also a random number $P_{N}\in\{2,3\ldots,N\}$ representing the number of potential parents at a timestep.
Let $\big\{(K^{(k)}_{N},\,P^{(k)}_{N})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random vectors that have the same distribution as $(K_{N},\,P_{N})$. For each $k\in\mathbb{Z}_{+}$, as before $K^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $k$-th time-step to be children. Then, suppose that (i) $P^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $(k+1)$-th time-step to be potential parents of the $K^{(k)}_{N}$ children, and (ii) the actual parent(s) of the child is (are) chosen uniformly without replacement from the $P^{(k)}_{N}$ potential parent(s).
@@
The model of NWF (25) corresponds to the special case when $K_{N}=1$ deterministically. When $\alpha_{N}=\alpha\in[0,1)$ for all $N$, this corresponds to the diploid Moran model with selfing considered in Lin (09). When $\alpha=0$ this corresponds to the model as in CLJ (22). The model in BBE (13) corresponds to our case when $\alpha_{N}=0$, and when $(K_{N},P_{N})$ is equal to $(1,2)$ and $([\psi N],2)$ with probabilities $1-\varepsilon_{N}$ and $\varepsilon_{N}$ respectively. There is a slight difference between these models and our model here, as the indices of the potential parents and those of the offspring are enforced to be disjoint in these models.
To simplify notation, we shall omit the subscript $N$ when there is no confusion. For example, we will write $(\alpha,K,S,K^{(k)})$ instead of $(\alpha_{N},K_{N},S_{N},K^{(k)}_{N})$.
Let $V_{i}:=\sum_{j=1}^{N}V_{i,j}$ is the total number of offspring for individual $i$, and $\tilde{V}_{i}:=V_{i,i}+V_{i}$. Then $\sum_{i=1}^{N}\tilde{V}_{i}=2K$ and the total genetic contribution of the $i$th individual to the next time step is
@@
In the absence of selfing ($\alpha_{N}=0$), we have $\tilde{V}_{i}=V_{i}$, recovering the offspring-frequency weights of MS (03) (see also BLS (18); ABFW (25)). These frequencies provide a convenient device to express one-step coalescence probabilities (e.g. pairwise and triple collisions) as explicit polynomials in $\{\tilde{V}_{i}\}$ with Mendelian coefficients.

### v2

@@ Abstract
We introduce a general diploid population model with self-fertilization (or ‘selfing’) and possible overlapping generations, and study the genealogy of a sample of $n$ genes as the population size $N$ tends to infinity. Unlike the traditional approach in coalescent theory which considers the unconditional (annealed) law of the gene genealogies averaged over the population pedigree, here we study the conditional (quenched) law of gene genealogies given the pedigree. We focus on the case of high selfing probability and obtain that this conditional law converges to a random probability measure. This measure is the random law of a system of coalescing random walks on an exchangeable fragmentation-coalescence process of Ber (04) which in general allows multiple mergers of ancestral lineages. As a special case, it contains a system of coalescing random walks on an ancestral graph with binary mergers only, which is identical in structure to the previously described ancestral recombination graph and ancestral selection graph. We use simulations of the ancestral process to show how the site-frequency spectrum of genetic data can depend strongly on the pedigree. Our convergence result is established by means of a characterization of weak convergence in distribution for random probability measures on Skorokhod spaces.
@@ 1 Introduction
This conceptual flaw was not unrecognized BNA (90); WKLR (12); WKW (16); WBLW (17); Ral (19), but perhaps luckily, it turned out that the Kingman coalescent can still be applied under the standard assumptions of neutral coalescent model because then the conditional limit is still equal to the unconditional limit Tyu (15). However, when there are simultaneous multiple mergers, the conditional genealogy can be different from the unconditional genealogy DFBW (24); ABFW (25).
@@
In ABFW (25), the authors study the diploid Cannings model introduced in BLS (18), where selfing is excluded. This is a model in which the offspring distribution, described by the matrix $(V_{ij})$ of the number of offspring between all pairs of parents, is invariant under permutations of the parents. Under the same condition on the offspring distribution that guaranteed annealed convergence to the $\Xi$-coalescent holds, the authors in ABFW (25) showed that the genealogies conditioned on the pedigree converge to an inhomogeneous $(\Psi,c_{\rm pair})$-coalescent, where $\Psi$ is a Poisson point process on $0,\infty)\times(\Delta\setminus\{\bf 0\})$ with intensity $dt\otimes\frac{1}{\langle x,x\rangle}\Xi(dx)$. This limiting process consists of the independent superposition of a Kingman coalescent with rate $c_{\rm pair}=1-\Xi(\Delta\setminus\{\bf 0\})$ and the multiple merger events whose the timing and intensity are specified by $\Psi$. The result in ABFW (25) generalizes earlier work in Tyu (15) and DFBW (24).
@@
While the model in BLS (18); ABFW (25) is quite general and captures a wide range of reproductive variance, it has two limitations from a biological perspective. Firstly, it does not allow for selfing, which is an important evolutionary force found in taxa including eukaryotic microbes and marine invertebrates (SR, 17; YSH, 23) and common in plants (AG, 89; HBG, 17; TWL+, 23). Selfing probabilities have a bi-modal distribution among plant species, with fewer species having $s\in(0.2,0.8)$ and some species reaching as high as 0.99 SL (85); AG (89); VK (01); BF (05); GKE (05); SAAMT (20). Secondly, the model assumes non-overlapping generations, a simplification that excludes biologically realistic scenarios in which parents may come from different age cohorts. Overlapping generations are common in many natural populations and can introduce temporal correlations and ancestral dependencies not captured by discrete-generation models.
@@
In previous work NWF (25) we considered a diploid Moran model with a high selfing probability $\alpha_{N}$, and obtained quenched limits for the coalescence time of a sample of size $n=2$. Conditioning coalescence on the pedigree revealed three markedly different behaviors, depending on how quickly $\alpha_{N}\to 1$. The critical case, called ‘limited outcrossing’ in NWF (25), is when $1-\alpha_{N}$ is of order $1/N$. This will also be our focus here.
@@
This paper aims to contribute to building the emerging quenched-coalescent theory, by going beyond the existing mathematical work Tyu (15); DFBW (24); NWF (25); ABFW (25). We introduce a general diploid model with overlapping generations that extends the haploid model of SW (08) and the diploid models of BLS (18) and NWF (25). Our model explicitly incorporates a selfing rate parameter and includes the diploid Moran-type models with selfing in NWF (25); CLJ (22); Lin (09) as special cases. Our main result significantly strengthens the analysis in NWF (25), in the ‘limited outcrossing’ regime, by considering arbitrary sample size $n$ and a substantially broader class of models. Moreover, it complements the results of ABFW (25); Tyu (15); DFBW (24) by identifying a distinct class of limiting conditional coalescent processes that arises in the high selfing regime.
@@
The new class of coalescent models that arise as $N\to\infty$ is a family of coalescing random walks on a directed random graph, which we call a $Q$-$\lambda$ graph in this paper. This random graph is a subgraph of an exchangeable fragmentation-coalescence process (EFC) introduced in Ber (04), where $Q$ is the rate matrix for coalescence and $\lambda$ is the fragmentation rate at which each node spits into two. The parameter $\lambda$ is the relative rate of outcrossing to pairwise coalescence as $N\to\infty$. In the regime where the classical (annealed) limit is a time-rescaled Kingman, we find that these $Q$-$\lambda$ graphs correspond to the ancestral graphs with binary mergers that have been used to describe genetic ancestries in the presence of recombination Hud (83); GM (97) and selection KN (97). As in NWF (25), time is rescaled so that any given pair of sample lineages coalesce with rate $2$ in the annealed limit in this paper. This is slightly different from that in ABFW (25) (where a given pair coalesce with rate 1) and that in DFBW (24) (which involves the parameter $\psi$); see Remarks 3.12 and 3.13 for detail. The proof of the main result (Theorem 4.7) involves a novel characterization of weak convergence in distribution of random probability measures on the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for any locally compact Polish space $E$, and a general theorem about quenched convergence of Markov processes on a suitably enriched space of partitions. While tightness and convergence of (deterministic) probability measures on Skorohod spaces have been studied (see BK (10); Kou (16)) by means of convergence determining and separating classes of functions, convergence of random probability measures on Skorohod spaces are less explored.
@@
From the application standpoint, the main motivation of coalescent theory (and our quenched-coalescent theory) is to describe the patterns of genetic variation expected under various biological scenarios to provide frameworks for statistical inference about past events and processes affecting populations. In a given population, gene genealogies are tree structures which emerge from tracing the ancestral lines of these sample backwards in time until the most recent common ancestor (MRCA). Mutations occurred in the past, along the ancestral lines, result in genetic diversity among the sample. For example, the site-frequency spectrum (SFS) is a commonly used measure of genetic variation upon which statistical inferences are based BWSH (01); Ach (09); EBBF (15); LF (15); GK (16); FLW+ (17); FKM+ (23). For a sample of $n$ genomes, the SFS records the number of polymorphic sites where a mutant base is found in $r\in\{1,\ldots,n-1\}$ copies Taj (89); BHK+ (95); Fu (95). Because per-site mutation rates are typically very small and the number of sites is large, the SFS is taken to reflect the total length of branches in the gene genealogy with $r$ descendants in the sample, or which are ancestral to $r$ of the samples. See Figure 5 for an illustration of these branches when $r=3$ and $n=10$. Different biological phenomena, such as population growth and natural selection, lead to different coalescent predictions for the SFS. Our main result sheds new light on the SFS of genetic data, by specifying how SFS depends on the pedigree. More precisely, under the conditional coalescent, the SFS should be viewed as a conditional SFS given the single pedigree as a latent variable. In Figure 6, we provide simulations for the conditional SFS given 5 different pedigrees and thus illustrate how the pedigree can impact the SFS of the data.
@@
Organization of this paper. Section 2 introduces our diploid exchangeable model with selfing and overlapping generations. An annealed version of our main result, akin to classical results of coalescent theory such as those of Kin (82); Möh (98); Pit (99); Sag (99); Sch (00), is presented in Section 3. Section 4 presents the main quenched convergence result for the coalescent of our exchangeable model model conditional on the pedigree. In Section 5 we present applications of our main convergence result and illustrate how the SFS can depend strongly on the pedigree. The remainder of the paper focuses on the tools and calculations necessary to prove the main results: weak convergence of random measures in distribution in Section 6, and a suitable notion of convergence of the random pedigree and continuity of the coalescent law’s dependence thereon in Section 7.
@@ 2.1 The model
In this paper, we introduce a diploid, monoecious, panmictic (well-mixed and randomly mating) population of constant size $N$, evolving in discrete time-steps with overlapping generations. This model generalizes the diploid model in ABFW (25); BLS (18) to allow overlapping generations and selfing, and extends the haploid model in SW (08) to diploid populations while supporting general offspring distributions.
@@
We consider discrete time-steps indexed by $k\in\mathbb{Z}_{+}=\{0,1,2,\ldots\}$, where $k=0$ is the present, $k=1$ the previous time-step, and so on backward into the past. Let $\big\{(K^{(k)}_{N},\textbf{V}^{(k)})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random variables that have the same distribution as $(K_{N},\textbf{V})$. Reproduction events in different time-steps are taken to be independent, and occur as follows:
For each $k\in\mathbb{Z}_{+}$, $K^{(k)}_{N}$ individuals are chosen uniformly without replacement, among the $N$ individuals from the $k$-th time-step, to be offspring. To describe parentage, we suppose that the individuals at each time-step are labeled by $[N]$. Let $V^{(k)}_{i,j}$ be the number of offspring in the $k$-th time step in the past, produced by the pair of individuals $(i,j)$ in the $(k+1)$-th time-step in the past. The $K_{N}^{(k)}$ individuals defined by the entries of $\textbf{V}^{(k)}$ constitute the new individuals in time-step $k$. Each individual has two sets of chromosomes, as it is a diploid population. Genetic lineages at an autosomal locus are transmitted according to Mendel’s law of random segregation, which means each gene copy in the offspring chooses independently from the two gene copies in the parent from which the gene copy is inherited.
@@
We further assume that each of the $K^{(k)}_{N}$ offspring independently chooses to have either a single parent (selfing) or two distinct parents (outcrossing), with probabilities $\alpha_{N}$ and $1-\alpha_{N}$ respectively. Therefore, conditional on $K_{N}^{(k)}$, the number of selfed offspring $\sum_{i=1}^{N}V_{i,i}^{(k)}$ is an independent binomial random variable: $\mathrm{Bin}(K_{N}^{(k)},\alpha_{N})$. In other words, we assume that $(K_{N},\textbf{V})$ satisfies
@@
Figure 1: An illustration of our population model between time-steps $k+1$ and $k$ in the past, with size $N=6$. There are $K_{N}=3$ offspring which are individuals 3,4,6 (from left to right) in timestep $k$. Bold edges represent reproductive relationships. Further, two of the offspring are reproduced by selfing. Individuals $1,2,5$, whose edges are marked in orange, simply persisted between consecutive time-steps and are not offspring. One can read $(V_{i,j})$ from the pedigree. We see that $V_{1}=0$ as the first parent (the first individual in timestep $k+1$) has no offspring, $V_{2}=2$ as the second parent has two offspring (individuals 3 and 4 in timestep $k$). Also, $V_{2,2}=V_{4,4}=1$.
@@
The population model described above includes two related but distinct classes of models in the literature as special cases.
@@
When $K_{N}=N$ (non-overlapping generations) and $\alpha_{N}=0$ (no selfing), our model reduces to the model introduced in BLS (18) and studied in ABFW (25).
@@
Our model also extends the model in SW (08) by modeling diploid, not just haploid, individuals and by allowing a more general offspring distribution.
Suppose the random symmetric matrices $\{V^{(k)}\}_{k\in\mathbb{Z}_{+}}$ satisfy the following extra assumption. Namely, their distribution is parametrized by a random number $P_{N}\in\{2,3\ldots,N\}$ representing the number of potential parents at a timestep.
Let $\big\{(K^{(k)}_{N},\,P^{(k)}_{N})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random vectors that have the same distribution as $(K_{N},\,P_{N})$. For each $k\in\mathbb{Z}_{+}$, as before $K^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $k$-th time-step to be offspring. Then, suppose that
  1. (i)
    $P^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $(k+1)$-th time-step to be potential parents of the $K^{(k)}_{N}$ offspring, and
  2. (ii)
    the actual parent(s) of the offspring is (are) chosen uniformly without replacement from the $P^{(k)}_{N}$ potential parent(s).
@@
The model of NWF (25) corresponds to the special case when $K_{N}=1$ deterministically. When $\alpha_{N}=\alpha\in[0,1)$ for all $N$, this corresponds to the diploid Moran model with selfing considered in Lin (09). When $\alpha=0$ this corresponds to the model as in CLJ (22). The model in BBE (13) corresponds to our case when $\alpha_{N}=0$, and when $(K_{N},P_{N})$ is equal to $(1,2)$ and $([\psi N],2)$ with probabilities $1-\varepsilon_{N}$ and $\varepsilon_{N}$ respectively. There is a slight difference between our model and these previous models, which assume that the indices of the potential parents and those of the offspring are disjoint.
To simplify notation, we shall omit the subscript $N$ when there is no confusion. For example, we will write $(\alpha,K,K^{(k)})$ instead of $(\alpha_{N},K_{N},K^{(k)}_{N})$.
Let $V_{i}:=\sum_{j=1}^{N}V_{i,j}$ be the total number of offspring for individual $i$, and $\tilde{V}_{i}:=V_{i,i}+V_{i}$. Then $\sum_{i=1}^{N}\tilde{V}_{i}=2K$ and the total genetic contribution of the $i$th individual to the next time step is
@@
In the absence of selfing ($\alpha_{N}=0$), we have $\tilde{V}_{i}=V_{i}$, recovering the offspring-frequency weights of MS (03) (see also BLS (18); ABFW (25)). These frequencies provide a convenient device to express one-step coalescence probabilities (e.g. pairwise and triple mergers) as explicit polynomials in $\{\tilde{V}_{i}\}$ with Mendelian coefficients.